# Size Generalization of Graph Neural Networks on Biological Data: Insights and Practices from the Spectral Perspective

### v4

@@ Appendix B Datasets
In this section, we provide additional details about the datasets utilized in our study. We utilize five pre-processed biological datasets, namely BBBP and BACE from the Open Graph Benchmark (Hu et al., 2020b), and PROTEINS, NCI1, and NCI109 from TuDataset (Morris et al., 2020), for graph classification in our experiments. Each graph in BBBP and BACE represents a molecule, where nodes are atoms, and edges are chemical bonds. Each node has a 9-dimensional feature vector, which contains its atomic number and chirality, as well as other additional atom features such as formal charge and whether the atom is in the ring or not (Hu et al., 2020b). In the PROTEINS dataset, nodes correspond to amino acids, and an edge connects two nodes if their distance is less than 6 Angstroms. Each node is associated with a 3-dimensional feature vector representing the type of secondary structure elements, such as helix, sheet, or turn. Each graph in NCI1 and NCI109 represents a chemical compound: each node stands for an atom and has a one-hot encoded feature representing the corresponding atom type; each edge represents the chemical bonds between the atoms. The description of each dataset is summarized as follows:
@@
  • NCI1 & NCI109: The NCI1 and NCI109 are two balanced subsets of chemical compounds screened for their activity against non-small cell lung cancer and ovarian cancer cell lines, respectively (Wale & Karypis, 2006). NCI1 contains a total of 4110 samples and NCI109 contains a total of 4117 samples.
@@ Appendix C Data Splits and Pre-processing
Table 5: Label distributions of BBBP, BACE, PROTEINS, NCI1 and NCI109 in the entire dataset, the smallest 50% subset, and the largest 10% subset.
@@
Table 6: Label distributions after proper data splits. both the Both the small_test and large_test subsets now have the same class distribution.
@@
Appendix D Additional Cycle-aware Baselines
In this section, we introduce the two additional baselines that are used to evaluate the effectiveness of cycle information in enhancing size generalization: +SSL and +AugCyc, which we briefly introduce in Section 5.
D.1 Self-supervised Auxiliary Task (+SSL)
This approach utilizes a self-supervised auxiliary task to enhance the node representations with cycle-related information. The auxiliary task is to predict whether a node belongs to a cycle. We do not utilize cycle lengths as labels because large test graphs may have cycle lengths not present in the training data. Formally, let $X^{(\text{Last})}$ denote the node representations obtained after the last graph convolution. The conventional way of learning is to directly apply a pooling operator and then minimize the loss function for label supervision as below:
$\mathcal{L}_{\text{label}}=\texttt{CrossEntropy}(\texttt{Linear}(\texttt{Pooling}({\mathbf{X}}^{(\text{Last})}),y^{{\mathcal{G}}}))$
$\mathcal{L}_{\text{cycle}}=\texttt{CrossEntropy}(\texttt{MLP}({\mathbf{X}}^{(\text{Last})}),{\mathbf{y}}_{\text{cycle}})$
$\mathcal{L}=\mathcal{L}_{\text{label}}+\lambda\mathcal{L}_{\text{cycle}}$
where $\lambda$ is a hyperparameter tuned via cross-validation.
D.2 Augmentation (+AugCyc)
This approach aims to reduce the discrepancy between small and large graphs through direct augmentation for the small training graphs. We augment the training graphs by extending the cycle lengths such that the average cycle length and standard deviation align with those in large graphs. To achieve this augmentation, we use the algorithm detailed in Section E.1. Additionally, we replicate the features from the nodes with the lowest degrees in the same cycle to populate the features for the newly added nodes. Last, we feed the augmented training graphs to GNNs for graph classification.
Appendix E Details and Discussions on Breaking Cycles and Aligning Cycle Lengths
E.1 Implementation Details
Our cycle-breaking algorithm is designed to remove all cycles in the graph with minimal edge removal while maintaining the current number of disconnected components. This approach minimally perturbs the graph while focusing on cycles. To achieve this, we begin by obtaining the cycle basis $\mathcal{C}$ (Section 2) of the input graph $\mathcal{G}$. Breaking all cycles can be accomplished by removing just one edge from each cycle in the basis. To ensure that this edge removal process does not increase the number of disconnected components, we propose a backtracking algorithm ( Algorithm 1). We note that while it is possible for Algorithm 1 to have failure cases, luckily, in all the datasets we have examined, we have not observed any instances where cycle breaking fails (i.e., Algorithm 1 returning False).
Algorithm 1 Cycle Breaking
[…]
2:     if  $i==\text{len}(\mathcal{C})$  then
3:         return True
5:     for  edge pair $\varepsilon$ in $\mathcal{C}_{i}$  do
6:          $\mathcal{G}\leftarrow\text{Remove}(\varepsilon)$
8:         if $N_{\text{before}}==N_{\text{after}}$ then
9:              if CycleBacktracking($i+1,\mathcal{G},\mathcal{C}$then
10:                  return True
11:         $\mathcal{G}\leftarrow\text{add}(\varepsilon)$
12:     return False
Table 7: Average cycle lengths for small and large graphs before and after cycle length alignment. Small and large graphs in real biological datasets originally differ in cycle length and such difference is reduced after cycle length alignment.
BBBP BACE NCI1 NCI109 PROTEINS
small large small large small large small large small large
Our cycle length alignment algorithm aligns both the mean and std of average cycle lengths between small and large graphs. This is achieved by selectively increasing the cycle length of a subset of small graphs. Formally, we present our algorithm in Algorithm 3. It has two hyperparameters: $n$, which controls the maximum cycle length increments one graph can get, and $\mathcal{R}$, which controls the portion of small graphs whose cycle lengths will be increased. We tune these hyperparameters in order to best align the cycle length of small graphs with that of large graphs. We tune $n$ over $[2,3,4,5,6,7]$ and $\mathcal{R}$ over $[1,2,3,4,5,6,7,8]$, and identify $n=7,\mathcal{R}=8$ for BBBP, $n=6,\mathcal{R}=8$ for BACE, $n=6,\mathcal{R}=5$ for NCI1, $n=6,\mathcal{R}=5$ for NCI109 and $n=5,\mathcal{R}=3$ for PROTEINS. We present the average cycle length after alignment in Table 7 and the discrepancy between small and large graphs is significantly reduced.
Algorithm 2 Add One Cycle Length
[…]
2:     for  each cycle $\text{cycle }\mathcal{C}_{i}$  do
3:         edge ($\text{edge }(v_{1},v_{2})\leftarrow\texttt{RandomChoice}(\mathcal{C}_{i})$
4:         $\mathcal{G}\leftarrow\text{RemoveEdge}(v_{1},v_{2})$
6:         $\mathcal{G}\leftarrow\text{AddNode}(v_{\text{new}})$
8:     return $\mathcal{G}$
Algorithm 3 Align Cycle Length
1:procedure AlignCycleLength($\{\mathcal{G}_{i}\}_{i=1}^{N}\colon\text{a set of graphs}$, :skipping $\mathcal{R}\colon\text{skipping ratio}$, $n\colon\text{increased cycle length},n\geq 1$)
2:     ResultList []$\text{ResultList }\mathcal{M}\leftarrow[]$
3:     for  $i=0,1,\dots,N-1$  do
4:         $\mathcal{G^{\prime}}\leftarrow\mathcal{G}_{i}$
5:         if $i\mod\mathcal{R}==0$ then
6:              for $j=0,1,\dots,n-1$ do
7:                  obtain Cycle Basis $\text{obtain Cycle Basis }\mathcal{C}\leftarrow\mathcal{G^{\prime}}$
9:         append $\mathcal{M}\leftarrow\text{append }\mathcal{G^{\prime}}$
10:     return $\mathcal{M}$
E.2 Ablation Study on Randomly Adding Nodes
One natural question arising from Section 3.3.1 is whether the cycle length alignment is necessary or it suffices to add nodes / edges randomly in order to reduce the spectrum discrepancy between small and large graphs. To answer this question, we conduct an ablation study on adding nodes and edges randomly to the graphs in the small dataset, matching the quantity introduced during the cycle length alignment process. We then compare the changes in the relative spectrum difference with those observed during the cycle length alignment procedure.
We present our results in Table 8. We find that aligning cycle lengths consistently leads to smaller spectrum differences between graphs of varying sizes compared to randomly adding nodes. These results further highlight the importance of cycle lengths for GNNs to achieve size generalizability.
Table 8: Comparison between randomly adding nodes and aligning cycle lengths. Average Wasserstein distance of eigenvalue distributions between graphs of similar size and different sizes are computed. Relative difference is computed as in Table 1. We use $\uparrow$ ($\downarrow$) to denote the increase (decrease) in the relative difference compared to not taking the corresponding action. Aligning cycle lengths results in a greater reduction of the relative spectrum difference compared to randomly adding a node.
Randonly adding nodes Aligning cycle lengths
BBBP 0.00557 0.00203 $\downarrow$ 33% 0.00565 0.00211 $\downarrow$ 41%
BACE 0.00406 0.00162 $\downarrow$ 26% 0.00417 0.00176 $\downarrow$ 41%
NCI1 0.00559 0.00230 $\downarrow$ 20% 0.00566 0.00242 $\downarrow$ 31%
@@ Appendix F Ablation Studies on Attention Mechanism
Table 9: Size generalizability evaluated with +NA, following the same rule as in Table 3. Naive attention is not helpful for size generalization.
@@
The results in Table 9 demonstrate that +NA does not consistently enhance graph classification for large graphs. This suggests that relying solely on the attention mechanism, without considering cycle information, does not improve the size generalizability of GNNs. In contrast, the incorporation of cycle information in +SIA leads to significant improvements on large graphs.
Appendix G Training Details
To ensure consistency with the original papers, we set the other hyperparameters of each baseline model to match the values specified in their respective papers. For +SSR, we use the same hyperparameter settings as in (Buffelli et al., 2022) and set the regularization coefficient $\lambda=0.1$ and coarsening ratios $C=\{0.8,0.9\}$ for all datasets. We adhere to the authors’ recommendation and employ spectral graph coarsening as the coarsening algorithm. Additionally, we fine-tune the aggregation strategy, selecting from options such as max, mean, and sum, to derive features for the nodes in the coarsened version. For SMP and +RPGNN, we use the original code provided by the authors. For SMP, we tune the number of SMP and FastSMP layers from the set {3,4}. Based on the validation results, we implement 3 FastSMP layers. For +RPGNN, we use the same hyperparameter settings as in (Bevilacqua et al., 2021; Buffelli et al., 2022). For +CIGA, we follow the same hyperparameter search range as specified in (Chen et al., 2022). Regarding the +SSL method that we proposed, we tune $\lambda$ over $\{1,1e-1,1e-2,1e-3,1e-4,1e-5\}$. Based on the validation results, we identify $\lambda=1e-2$ for NCI1, and NCI109, $\lambda=1$ for PROTEINS, BBBP, and $\lambda=1e-2$ for BACE. For our proposed cycle augmentation, we use the set of hyperparameters $n,\mathcal{R}$ from Section E.1. For +SIA, we have no hyperparameters to tune.
Appendix H Additional Plots on Degree Distributions
Figure 3 and Table 10 reveal that the degree distribution does not exhibit a clear correlation with the graph size. These visualizations and tables adhere to the same convention as presented in Section 3.2. Unlike the assumption made in (Yehudai et al., 2021) that attributes changes in local degree patterns as the primary covariate shifts induced by size, our investigation on biological datasets suggests a different perspective.
Figure 3: Pairwise graph distance measured by the Wasserstein distance of degree distributions, following the same convention in Figure 1. Degree distribution does not show a clear correlation with graph size.
Table 10: Average Wasserstein distance between graphs of ‘similar sizes’ and graphs of ‘different sizes’ based on degree distributions, respectively. The relative difference is computed by the difference of the Wasserstein distance normalized by the Wasserstein distance of similar graphs.
BBBP BACE PROTEINS NCI109 NCI1
Different Size 0.00177 0.00120 0.00062 0.00161 0.00163
Similar Size 0.00157 0.00105 0.00054 0.00146 0.00148
Relative Difference 12.1% 14.6% 14.8% 10.1% 9.8%
Appendix I Additional Results on Real-world Datasets
In this section, we demonstrate +SIA’s effectiveness in size generalization using the HIV dataset from (Hu et al., 2020a). Adopting the ROC-AUC metric from (Chen et al., 2022; Hu et al., 2020a), we observe that +SIA still consistently achieves performance improvements across all backbone models.
Table 11: Size generalizability evaluated on the real-world hiv datasets. +SIA still leads to consistent performance boosts.
Dataset Hiv
Models MLP GCN GAT GIN FAGCN GNNML3
Small Large Small Large Small Large Small Large Small Large Small Large
Original Performance 67.95$\pm$2.22 38.97$\pm$6.39 71.29$\pm$1.58 38.35$\pm$4.09 73.34$\pm$1.77 41.23$\pm$4.62 71.04$\pm$1.45 43.65$\pm$3.01 68.35$\pm$3.09 50.03$\pm$1.85 74.32$\pm$1.13 47.44$\pm$7.66
+SIA Performance 70.64$\pm$1.94 40.93$\pm$2.87 70.27$\pm$1.42 43.24$\pm$6.08 72.33$\pm$1.72 46.72$\pm$2.74 69.92$\pm$1.50 43.55$\pm$3.84 69.53$\pm$1.92 52.35$\pm$6.35 74.52$\pm$2.95 54.68$\pm$6.60
+Avg Improvements (small) +0.15 ; (large) +3.63
Appendix J Additional Plots on Cycle Breaking & Cycle Length Alignment
@@ Appendix I Additional Results on Real-world Datasets
BBBP
BACE
PROTEINS
NCI1
Figure 4: Pairwise graph distance measured by the Wasserstein distance of eigenvalue distributions after breaking cycles and aligning cycle lengths on different datasets. Breaking cycles amplifies the correlation between eigenvalue distribution and graph size, while aligning cycle lengths reduces the correlation.
Figure 5: A simple counter-example of d-pattern. The d-pattern for each pair of nodes with the same id (number) in figures (a) and (b) is identical (proved in K.3). The figure follows the same convention as Figure 2 of Yehudai et al. (2021), where the left side represents the two non-isomorphic graphs with 7 nodes and each color represents a different feature. The right side illustrates the 0, 1, and 2-pattern of the node with id 2.
Figure 6: Illustration of recursive updates of node 2 in (a) and (b) from Figure 5. We can find that in each hierarchy the d-pattern is the same.

### v5

@@ Appendix B Datasets
In this section, we provide additional details about the datasets utilized in our study. We utilize five pre-processed biological datasets, namely BBBP and BACE from the Open Graph Benchmark (Hu et al., 2020), and PROTEINS, NCI1, and NCI109 from TuDataset (Morris et al., 2020), for graph classification in our experiments. Each graph in BBBP and BACE represents a molecule, where nodes are atoms, and edges are chemical bonds. Each node has a 9-dimensional feature vector, which contains its atomic number and chirality, as well as other additional atom features such as formal charge and whether the atom is in the ring or not (Hu et al., 2020). In the PROTEINS dataset, nodes correspond to amino acids, and an edge connects two nodes if their distance is less than 6 Angstroms. Each node is associated with a 3-dimensional feature vector representing the type of secondary structure elements, such as helix, sheet, or turn. Each graph in NCI1 and NCI109 represents a chemical compound: each node stands for an atom and has a one-hot encoded feature representing the corresponding atom type; each edge represents the chemical bonds between the atoms. The description of each dataset is summarized as follows:
@@
  • NCI1 & NCI109: The NCI1 and NCI109 are two balanced subsets of chemical compounds screened for their activity against non-small cell lung cancer and ovarian cancer cell lines, respectively (Wale and Karypis, 2006). NCI1 contains a total of 4110 samples and NCI109 contains a total of 4117 samples.
@@ Appendix C Data Splits and Pre-processing
Table 8. Label distributions of BBBP, BACE, PROTEINS, NCI1 and NCI109 in the entire dataset, the smallest 50% subset, and the largest 10% subset.
@@
Table 9. Label distributions after proper data splits. both the Both the small_test and large_test subsets now have the same class distribution.
@@
Appendix D Training Details
To ensure consistency with the original papers, we set the other hyperparameters of each baseline model to match the values specified in their respective papers. For +SSR, we use the same hyperparameter settings as in (Buffelli et al., 2022) and set the regularization coefficient $\lambda=0.1$ and coarsening ratios $C=\{0.8,0.9\}$ for all datasets. We adhere to the authors’ recommendation and employ spectral graph coarsening as the coarsening algorithm. Additionally, we fine-tune the aggregation strategy, selecting from options such as max, mean, and sum, to derive features for the nodes in the coarsened version. For SMP and +RPGNN, we use the original code provided by the authors. For SMP, we tune the number of SMP and FastSMP layers from the set {3,4}. Based on the validation results, we implement 3 FastSMP layers. For +RPGNN, we use the same hyperparameter settings as in (Bevilacqua et al., 2021; Buffelli et al., 2022). For +CIGA, we follow the same hyperparameter search range as specified in (Chen et al., 2022). Regarding the +SSL method that we proposed, we tune $\lambda$ over $\{1,1e-1,1e-2,1e-3,1e-4,1e-5\}$. Based on the validation results, we identify $\lambda=1e-2$ for NCI1, and NCI109, $\lambda=1$ for PROTEINS, BBBP, and $\lambda=1e-2$ for BACE. For our proposed cycle augmentation, we use the set of hyperparameters $n,\mathcal{R}$ from Section G.1. For +SIA, we have no hyperparameters.
Appendix E Additional Tables
To facilitate the interpretation of the experimental results, we present the averaged improvements on both small and large graphs, evaluated across various model backbones and datasets, in Table 11. Our findings suggest that incorporating cycle information enhances size generalization, with +SIA demonstrating the most significant improvements. In addition, we show the per-epoch runtime in Table 10.
Table 10. Comparison of per-epoch time across different methods, measured in minutes.
bbbp bace proteins NCI1 NCI109
Original 0.33 0.21 0.13 0.55 0.53
+SSR 1.55 0.93 0.74 2.36 2.22
+CIGAv2 3.10 1.92 1.32 4.25 4.11
+SSL (Ours) 0.33 0.21 0.12 0.57 0.54
+AugCyc (Ours) 0.34 0.21 0.13 0.54 0.53
+SIA (Ours) 1.32 0.78 0.55 1.98 1.94
Table 11. Averaged experimental results for model-agnostic methods evaluated across various model backbones and datasets. The results indicate that +SIA achieves the most significant overall improvements.
Method SIA SSL AUG GRPGNN SSR CIGAv1 CIGAv2
Avg Improv (Small) 1.09 0.78 0.07 0.37 0.73 0.24 -0.05
Avg Improv (Large) 6.02 2.55 3.44 1.15 2.43 3.24 4.29
@@ Appendix F Ablation Studies on Attention Mechanism
Table 12. Size generalizability evaluated with +NA, following the same rule as in Table 3. Naive attention is not helpful for size generalization.
@@
We use +NA to denote the use of naive attention without using cycle features. The results in Table 12 demonstrate that +NA does not consistently enhance graph classification for large graphs. This suggests that relying solely on the attention mechanism, without considering cycle information, does not improve the size generalizability of GNNs. In contrast, the incorporation of cycle information in +SIA leads to significant improvements on large graphs.
Appendix G Details and Discussions on Breaking Cycles and Aligning Cycle Lengths
G.1. Implementation Details
Our cycle-breaking algorithm is designed to remove all cycles in the graph with minimal edge removal while maintaining the current number of disconnected components. This approach minimally perturbs the graph while focusing on cycles. To achieve this, we begin by obtaining the cycle basis $\mathcal{C}$ (Section 2) of the input graph $\mathcal{G}$. Breaking all cycles can be accomplished by removing just one edge from each cycle in the basis. To ensure that this edge removal process does not increase the number of disconnected components, we propose a backtracking algorithm ( Algorithm 1). We note that while it is possible for Algorithm 1 to have failure cases, luckily, in all the datasets we have examined, we have not observed any instances where cycle breaking fails (i.e., Alg. 1 returning False).
Algorithm 1 Cycle Breaking
[…]
2:  if  $i==\text{len}(\mathcal{C})$  then
3:   return True
5:  for  edge pair $\varepsilon$ in $\mathcal{C}_{i}$  do
6:    $\mathcal{G}\leftarrow\text{Remove}(\varepsilon)$
8:   if $N_{\text{before}}==N_{\text{after}}$ then
[…]
10:      return True
11:   $\mathcal{G}\leftarrow\text{add}(\varepsilon)$
12:  return False
Table 13. Average cycle lengths for small and large graphs before and after cycle length alignment. Small and large graphs in real biological datasets originally differ in cycle length and such difference is reduced after cycle length alignment.
BBBP BACE NCI1 NCI109 PROTEINS
small large small large small large small large small large
Our cycle length alignment algorithm aligns both the mean and std of average cycle lengths between small and large graphs. This is achieved by selectively increasing the cycle length of a subset of small graphs. Formally, we present our algorithm in Algorithm 3. It has two hyperparameters: $n$, which controls the maximum cycle length increments one graph can get, and $\mathcal{R}$, which controls the portion of small graphs whose cycle lengths will be increased. We tune these hyperparameters in order to best align the cycle length of small graphs with that of large graphs. We tune $n$ over [2,3,4,5,6,7][2,3,4,5,6,7][ 2 , 3 , 4 , 5 , 6 , 7 ] and $\mathcal{R}$ over [1,2,3,4,5,6,7,8][1,2,3,4,5,6,7,8][ 1 , 2 , 3 , 4 , 5 , 6 , 7 , 8 ], and identify $n=7,\mathcal{R}=8$ for BBBP, $n=6,\mathcal{R}=8$ for BACE, $n=6,\mathcal{R}=5$ for NCI1, $n=6,\mathcal{R}=5$ for NCI109 and $n=5,\mathcal{R}=3$ for PROTEINS. We present the average cycle length after alignment in Table 13 and the discrepancy between small and large graphs is significantly reduced.
Algorithm 2 Add One Cycle Length
[…]
2:  for  each $\text{cycle }\mathcal{C}_{i}$  do
3:   $\text{edge }(v_{1},v_{2})\leftarrow\texttt{RandomChoice}(\mathcal{C}_{i})$
4:   $\mathcal{G}\leftarrow\text{RemoveEdge}(v_{1},v_{2})$
6:   $\mathcal{G}\leftarrow\text{AddNode}(v_{\text{new}})$
8:  return $\mathcal{G}$
Algorithm 3 Align Cycle Length
[…]
2:  $\text{ResultList }\mathcal{M}\leftarrow[]$
3:  for  $i=0,1,\dots,N-1$  do
4:   $\mathcal{G^{\prime}}\leftarrow\mathcal{G}_{i}$
5:   if $i\mod\mathcal{R}==0$ then
6:     for $j=0,1,\dots,n-1$ do
7:      $\text{obtain Cycle Basis }\mathcal{C}\leftarrow\mathcal{G^{\prime}}$
9:   $\mathcal{M}\leftarrow\text{append }\mathcal{G^{\prime}}$
10:  return $\mathcal{M}$
G.2. Ablation Study on Randomly Adding Nodes
One natural question arising from Section 3.2.1 is whether the cycle length alignment is necessary or it suffices to add nodes / edges randomly in order to reduce the spectrum discrepancy between small and large graphs. To answer this question, we conduct an ablation study on adding nodes and edges randomly to the graphs in the small dataset, matching the quantity introduced during the cycle length alignment process. We then compare the changes in the relative spectral difference with those observed during the cycle length alignment procedure.
We present our results in Table 14. We find that aligning cycle lengths consistently leads to smaller spectral differences between graphs of varying sizes compared to randomly adding nodes. These results further highlight the importance of cycle lengths for GNNs to achieve size generalizability.
Table 14. Comparison between randomly adding nodes and aligning cycle lengths. Average Wasserstein distance of eigenvalue distributions between graphs of similar size and different sizes are computed. Relative difference is computed as in Table 1. We use $\uparrow$ ($\downarrow$) to denote the increase (decrease) in the relative difference compared to not taking the corresponding action. Aligning cycle lengths results in a greater reduction of the relative spectral difference compared to randomly adding a node.
Randonly adding nodes Aligning cycle lengths
BBBP 0.00557 0.00203 $\downarrow$ 33% 0.00565 0.00211 $\downarrow$ 41%
BACE 0.00406 0.00162 $\downarrow$ 26% 0.00417 0.00176 $\downarrow$ 41%
NCI1 0.00559 0.00230 $\downarrow$ 20% 0.00566 0.00242 $\downarrow$ 31%
Appendix H Additional Plots on Cycle Breaking & Cycle Length Alignment
@@
Appendix I Additional Plots on Scaled Eigenvector
Similar to Figure 1, Figure 3 presents the pairwise graph distance measured using scaled eigenvectors across the five datasets. We observe that the dependence of the scaled eigenvectors on graph size is mild.
Figure 3. Pairwise graph distance measured by the Wasserstein distance of the scaled eigenvector distributions, evaluated using the same metrics as in Figure 1. The scaled eigenvector does not show a clear correlation with graph size.
Appendix J Additional Plots on Degree Distributions
Figure 4. Pairwise graph distance measured by the Wasserstein distance of degree distributions, following the same convention in Figure 1. Degree distribution does not show a clear correlation with graph size.
Figure 4 reveal that the degree distribution does not exhibit a clear correlation with the graph size. These visualizations and tables adhere to the same convention as presented in Section 3.2. Unlike the assumption made in (Yehudai et al., 2021) that attributes changes in local degree patterns as the primary covariate shifts induced by size, our investigation on biological datasets suggests a different perspective.